# Electron-phonon coupling induced topological phase transition in an $\alpha$-$T_{3}$ Haldane-Holstein model

### v1

@@ III.4 Hall conductivity
Interestingly, for $\alpha=0.8$, the quantized (tiny) Hall plateaus in Fig. 16$(d)$ are located at $e^{2}/h$ (where, $|C|=1$) and $2e^{2}/h$ (where, $|C|=2$) for $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ (shown for $\lambda_{eph}=0.3$ and $0.4$) and $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regimes (shown for $\lambda_{eph}=0.1$, denoted by solid green), respectively, confirm existence of two distinct topological insulating phases with a nonzero $\sigma_{xy}$ (also true for any $\alpha$ in $10.75\lesssim\alpha<$ 1 regime), while these plateaus vanish beyond $\lambda_{c_{2}}$, ascertaining emergence of a trivial insulating phase ($C=0$) with $\sigma_{xy}=0$. Thus, the polaronic Hall conductivity ensures that the system undergoes a transition from a nontrivial insulating phase with quantized plateaus at $|C|e^{2}/h$ in the $\lambda_{eph}\lesssim\lambda_{c}$ regime to a trivial insulating phase with zero Hall conductivity in the $\lambda_{eph}>\lambda_{c}$ regime.
@@ IV Conclusion
To summarize, we have studied the effect of e-ph interaction on inducing a topological phase transition in a Haldane-Holstein model on an $\alpha$-$T_{3}$ lattice. The NN and the complex NNN Haldane hopping amplitudes get renormalized by the Holstein reduction factor showing the signature of polaron formation in the system. The cases of our study are majorly divided into two scenarios, namely lower to intermediate $\alpha$ $(0<\alpha\leq 0.6)$ and higher $(0.6<\alpha<1)$ values of $\alpha$. With the help of the effective Hamiltonian in $\bf{k}$-space, we have computed the bulk and the edge spectra where it is observed that for the first case, as we increase the e-ph coupling strength $\lambda_{eph}$, the bulk gap between the flat and valance bands closes at a critical coupling strength, namely $\lambda_{c}$ at one $\bf{K}$-valley and re-opens beyond $\lambda_{c}$. This feature explains that the system is characterized by two distinct insulating states below and above $\lambda_{eph}=\lambda_{c}$. Consequently, the conducting edge modes emerge in the $\lambda_{eph}<\lambda_{c}$ regime, which are preserved up to $\lambda_{eph}=\lambda_{c}$, and disappear for $\lambda_{eph}>\lambda_{c}$. In the second case, we encounter a different scenario where the flat and valance bands in the bulk remain gapless for the $0.65\lesssim\alpha\lesssim 0.75$ regime and gapped for $10.75\lesssim\alpha<$ 1, till $\lambda_{eph}$ reaches a first critical value, namely $\lambda_{c_{1}}$ and become gapped till $\lambda_{eph}$ assumes another critical value, namely $\lambda_{c_{2}}$ where similar gap closing transition takes place. The explicit emergence of conducting edge modes in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime both for $0.65\lesssim\alpha\lesssim 0.75$ and $10.75\lesssim\alpha<$ 1, and also in $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime for $10.75\lesssim\alpha<$ 1 that traverse through the flat band around the $\bf{K}$ and $\bf{K^{\prime}}$ valleys makes the latter case more intriguing. It indicates that for $0.65\lesssim\alpha\lesssim 0.75$ $(0.75\lesssim\alpha<1)$, the system re-enters from a conventional (topological) $SM$ phase (in the $\lambda_{eph}<\lambda_{c_{1}}$ regime) to a (another) topological one (in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime) upon tuning the e-ph coupling strength. The above discussions for both the cases, either with a unique $\lambda_{c}$ or with two $\lambda_{c}$s, namely $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$, strongly indicate possibilities of inducing topological phase transition via e-ph coupling in an $\alpha$-$T_{3}$ Haldane-Holstein model. Furthermore, we have numerically computed the Berry curvature and the topological invariant, namely the (polaronic) Chern number ($C$), for different values of $\alpha$. In our study, the evidence of a discontinuous change in $C$ from $|C|=1$ to $|C|=0$ for $0<\alpha\lesssim 0.75$ regime, and from $|C|=2$ to $|C|=1$ and finally to $|C|=0$ for $10.75\lesssim\alpha<$ 1 regime exhibiting a jump in the $C$ vs $\lambda_{eph}$ diagram at different critical values of the e-ph coupling for different values of $\alpha$ directly confirms the topological phase transition solely caused by the e-ph interaction, while interpolating $\alpha$ between corresponding lattice structures of graphene to a dice lattice. More specifically, the system under investigation possesses a topological insulating phase accompanied by $|C|=1$ or $|C|=2$ (depending on the range of $\alpha$) below certain critical values of the e-ph coupling strength, and becomes a trivial insulator ($C=0$) above the critical point. We, furthermore, incorporate the above observations in a phase diagram plotted for $C$ in the $\lambda_{eph}-\alpha$ plane. To confirm such phases, and phase transitions from one phase to another, we have calculated the Hall conductivity for a few values of $\alpha$ (both small and large) as a function of $\lambda_{eph}$. The existence of Hall plateaus at $|C|e^{2}/h$ below a certain critical $\lambda_{c}$ for a particular value of $\alpha$ further substantiates the evidence of topological phase transitions induced by e-ph coupling in our $\alpha$-$T_{3}$ Haldane-Holstein model. We wish to motivate that our study may serve as a powerful tool for understanding the interaction-driven topology in novel quantum systems.
@@ Acknowledgements
K.B. sincerely thanks IIT Guwahati for providing financial support through the Institute Post Doctoral Fellowship (Ref no. IITG/R&\&&D/IPDF/2023-24/20231003P537). M.I. and K.B. also sincerely acknowledge Ms. Srijata Lahiri and Dr. Sayan Mondal for fruitful discussions.
Appendix A Derivation of the modified Haldane-Holstein Hamiltonian for an ${\boldsymbol{\alpha\--T_{3}}}$ lattice
@@
$[R,[R,\mathcal{H}^{(12)}]]=-(1+\alpha)t\sum_{i,\delta}c^{\dagger}_{i}c_{i+\delta}\biggl{[}\lambda_{eph}[X_{i}-X_{i+\delta}]\biggr{]}^{2},~{}[R,[R,[R,\mathcal{H}^{(12)}]]]=-(1+\alpha)t\sum_{i,\delta}c^{\dagger}_{i}c_{i+\delta}\biggl{[}\lambda_{eph}[X_{i}-X_{i+\delta}]\biggr{]}^{3}.$ (19)
where $X_{i}\equiv b^{\dagger}_{i}-b_{i}$ and $\delta$ is the NN index, that is, $j=i+\delta$. Therefore, collecting terms in Eq. (18) and Eq. (19) and using Eq. (17) the NN terms are transformed as
~$\tilde{\mathcal{H}}^{(12)}=-(1+\alpha)t\sum_{i,\delta}c^{\dagger}_{i}c_{i+\delta}\biggl{[}1+\lambda_{eph}[X_{i}-X_{i+\delta}]+\frac{1}{2!}\biggl{[}\lambda_{eph}[X_{i}-X_{i+\delta}]\biggr{]}^{2}+\frac{1}{3!}\biggl{[}\lambda_{eph}[X_{i}-X_{i+\delta}]\biggr{]}^{3}+...\biggr{]}$
$=-(1+\alpha)t\sum_{i,\delta}c^{\dagger}_{i}c_{i+\delta}~{}e^{\lambda_{eph}[X_{i}-X_{i+\delta}]}.$ (20)
@@
~$\tilde{\mathcal{H}}^{(34)}=-(1+\alpha)\frac{i\lambda}{3\sqrt{3}}\sum_{i,\eta}c^{\dagger}_{i}c_{i+\eta}~{}e^{\lambda_{eph}[X_{i}-X_{i+\eta}]}e^{i\phi_{i,i+\eta}},$ (21)

### v2

@@ III.4 Hall conductivity
Interestingly, for $\alpha=0.8$, the quantized (tiny) Hall plateaus in Fig. 16$(d)$ are located at $e^{2}/h$ (where, $|C|=1$) and $2e^{2}/h$ (where, $|C|=2$) for $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ (shown for $\lambda_{eph}=0.3$ and $0.4$, denoted by solid blue and black, respectively) and $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regimes (shown for $\lambda_{eph}=0.1$, denoted by solid green), respectively, confirm existence of two distinct topological insulating phases with a nonzero $\sigma_{xy}$ (also true for any $\alpha$ in $10.75\lesssim\alpha<$ 1 regime), while these plateaus vanish beyond $\lambda_{c_{2}}$, ascertaining emergence of a trivial insulating phase ($C=0$) with $\sigma_{xy}=0$. Thus, the polaronic Hall conductivity ensures that the system undergoes a transition from a nontrivial insulating phase with quantized plateaus at $|C|e^{2}/h$ in the $\lambda_{eph}\lesssim\lambda_{c}$ regime to a trivial insulating phase with zero Hall conductivity in the $\lambda_{eph}>\lambda_{c}$ regime.
@@ IV Conclusion
To summarize, we have studied the effect of e-ph interaction on inducing a topological phase transition in a Haldane-Holstein model on an $\alpha$-$T_{3}$ lattice. The NN and the complex NNN Haldane hopping amplitudes get renormalized by the Holstein reduction factor showing the signature of polaron formation in the system. The cases of our study are majorly divided into two scenarios, namely lower to intermediate $\alpha$ $(0<\alpha\leq 0.6)$ and higher $(0.6<\alpha<1)$ values of $\alpha$. With the help of the effective Hamiltonian in $\bf{k}$-space, we have computed the bulk and the edge spectra where it is observed that for the first case, as we increase the e-ph coupling strength $\lambda_{eph}$, the bulk gap between the flat and valance bands closes at a critical coupling strength, namely $\lambda_{c}$ at one $\bf{K}$-valley and re-opens beyond $\lambda_{c}$. This feature explains that the system is characterized by two distinct insulating states below and above $\lambda_{eph}=\lambda_{c}$. Consequently, the conducting edge modes emerge in the $\lambda_{eph}<\lambda_{c}$ regime, which are preserved up to $\lambda_{eph}=\lambda_{c}$, and disappear for $\lambda_{eph}>\lambda_{c}$, signifying a topologically nontrivial to trivial phase transition. In the second case, we encounter a different scenario where the flat and valance bands in the bulk remain gapless for the $0.65\lesssim\alpha\lesssim 0.75$ regime and gapped for $10.75\lesssim\alpha<$ 1, till $\lambda_{eph}$ reaches a first critical value, namely $\lambda_{c_{1}}$ and become gapped till $\lambda_{eph}$ assumes another critical value, namely $\lambda_{c_{2}}$ where similar gap closing transition takes place. The explicit emergence of conducting edge modes in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime both for $0.65\lesssim\alpha\lesssim 0.75$ and $10.75\lesssim\alpha<$ 1, and also in $0\leq\lambda_{eph}<\lambda_{c_{1}}$ regime for $10.75\lesssim\alpha<$ 1 that traverse through the FB around the $\bf{K}$ and $\bf{K^{\prime}}$ valleys makes the latter case more intriguing. It indicates that for $0.65\lesssim\alpha\lesssim 0.75$ $(0.75\lesssim\alpha<1)$, the system re-enters from a conventional (topological) $SM$ phase (in the $\lambda_{eph}<\lambda_{c_{1}}$ regime) to a (another) topological one (in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime) upon tuning the e-ph coupling strength. The above discussions for both the cases, either with a unique $\lambda_{c}$ or with two $\lambda_{c}$s, namely $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$, strongly indicate possibilities of inducing topological phase transition via e-ph coupling in an $\alpha$-$T_{3}$ Haldane-Holstein model. Furthermore, we have numerically computed the Berry curvature and the topological invariant, namely the (polaronic) Chern number ($C$), for different values of $\alpha$. In our study, the evidence of a discontinuous change in $C$ from $|C|=1$ to $|C|=0$ for $0<\alpha\lesssim 0.75$ regime, and from $|C|=2$ to $|C|=1$ and finally to $|C|=0$ for $10.75\lesssim\alpha<$ 1 regime exhibiting a jump in the $C$ vs $\lambda_{eph}$ diagram at different critical values of the e-ph coupling for different values of $\alpha$ directly confirms the topological phase transition solely caused by the e-ph interaction, while interpolating $\alpha$ between corresponding lattice structures of graphene to a dice lattice. More specifically, the system under investigation possesses a topological insulating phase accompanied by $|C|=1$ or $|C|=2$ (depending on the range of $\alpha$) below certain critical values of the e-ph coupling strength, and becomes a trivial insulator ($C=0$) above the critical point. We, furthermore, incorporate the above observations in a phase diagram plotted for $C$ in the $\lambda_{eph}-\alpha$ plane. To confirm such phases, and phase transitions from one phase to another, we have calculated the Hall conductivity for a few values of $\alpha$ (both small and large) as a function of $\lambda_{eph}$. The existence (vanishing) of Hall plateaus at $|C|e^{2}/h$ below (above) a certain critical $\lambda_{c}$ for a particular value of $\alpha$ further substantiates the evidence of topological phase transitions induced by e-ph coupling in our $\alpha$-$T_{3}$ Haldane-Holstein model. We wish to motivate that our study may serve as a powerful tool for understanding the interaction-driven topology in novel quantum systems.
@@ Acknowledgements
K.B. sincerely thanks IIT Guwahati for providing financial support through the Institute Post Doctoral Fellowship (Ref. No. IITG/R&\&&D/IPDF/2023-24/20231003P537). M.I. and K.B. also sincerely acknowledge Ms. Srijata Lahiri and Dr. Sayan Mondal for fruitful discussions.
Appendix A Derivation of the modified Haldane-Holstein Hamiltonian for an $\alpha$-$T_{3}$ lattice
@@
$[R,[R,\mathcal{H}^{(12)}]]=-(1+\alpha)t\sum_{i,\delta}c^{\dagger}_{i}c_{i+\delta}[X_{i}-X_{i+\delta}]^{2},~{}[R,[R,[R,\mathcal{H}^{(12)}]]]=-(1+\alpha)t\sum_{i,\delta}c^{\dagger}_{i}c_{i+\delta}[X_{i}-X_{i+\delta}]^{3}.$ (19)
where $X_{i}\equiv\lambda_{eph}(b^{\dagger}_{i}-b_{i})$ and $\delta$ is the NN index, that is, $j=i+\delta$. Therefore, collecting terms in Eq. (18) and Eq. (19) and using Eq. (17) the NN terms are transformed as
~$\tilde{\mathcal{H}}^{(12)}=-(1+\alpha)t\sum_{i,\delta}c^{\dagger}_{i}c_{i+\delta}\biggl{[}1+[X_{i}-X_{i+\delta}]+\frac{1}{2!}[X_{i}-X_{i+\delta}]^{2}+\frac{1}{3!}[X_{i}-X_{i+\delta}]^{3}+...\biggr{]}$
$=-(1+\alpha)t\sum_{i,\delta}c^{\dagger}_{i}c_{i+\delta}~{}e^{[X_{i}-X_{i+\delta}]}.$ (20)
@@
~$\tilde{\mathcal{H}}^{(34)}=-(1+\alpha)\frac{\lambda}{3\sqrt{3}}\sum_{i,\eta}c^{\dagger}_{i}c_{i+\eta}~{}e^{[X_{i}-X_{i+\eta}]}e^{i\phi_{i,i+\eta}},$ (21)